Consider an example of a new car sale yard and the managers predicted that they would sell 1,400 of model A and 1,800 of model B in the fiscal year. They would prepare their budget accordingly. Sales of model A did not meet expectations and sold 1,200 units. Model B had sales rise to 2,400 vehicles for the year. The expected margin on each vehicle is $2,000 for model A and $3,000 for model B.

Calculate the predicted and actual sales mix and the impact of the variations on income.

please do not copy and paste from other answers as they are wrong
Answer #1

Answer:-

Budgeted sales quantity (BQ):

    Model A     1,400 units

    Model B     1,800 units

Actual sales quantity (AQ):

    Model A     1,200 units

    Model B    2,400 units

Revised actual sales quantity (RAQ):

    Model A = 1,400/(1,400 + 1,800) × (1,200 + 2,400) = 1,575

    Model B = 1,800/(1,400 + 1,800) × (1,200 + 2,400) = 2,025

Budgeted margin (BM):

    Model A     $2,000 per vehicle

    Model B     $3,000 per vehicle

Model AQ × BM RAQ × BM
A 1,200 × $2,000 1,575 × $2,000
B 2,400 × $3,000 2,025 × $3,000
Total $       9,600,000 $       9,225,000

Sales mix variance = (RAQ × BM) - (AQ × BM)

                               = $9,225,000 - $9,600,000 = 375,000 Adverse

Predicted sales mix = $9,225,000

Actual sales mix = $9,600,000

Analysis: It can be seen that there is an adverse of 375,000 which has arisen due increase in mix of model B which is also having higher margin than model A. This will increase income.